For what values of a and b is

g(x)= { ax +2b, x is less than and equal to -2 }
{ x squared + 3a- b, -2< x less than and equal to 3 }
{3x-5, x> 3}
1
Expert's answer
2019-08-12T11:38:38-0400

In order for the function to be continuous it is necessary that the pieces of function "g(x)" are equal in points of discontinuity.

The points of discontinuity are the x = -2 and x = 3. Then:

"\begin{cases}
 ax + 2b = x^2 +3a - b, x= -2 
 \\
 x^2 +3a - b = 3x - 5, x = 3
 \end{cases}" "\implies" "\begin{cases}
 -2a + 2b = 4 + 3a - b, 
 \\
 9 +3a - b = 9 - 5
 \end{cases}" "\implies" "\begin{cases}
 3b = 4 + 5a, 
 \\
 3a - b = -5
 \end{cases}""\implies"


"\begin{cases}
 b = \frac{4 + 5a}{3} ,
 \\
 b = 3a + 5
 \end{cases}" "\implies" "\frac{4 + 5a}{3} = 3a + 5"


"4 + 5a = 9a + 15 \implies a = -\frac{11}{4}" "\implies b = 3* (-\frac{11}{4}) + 5 = -\frac{13}{4}"


Therefore, "g(x)" is continuous when "a = -\frac{11}{4} , b = -\frac{13}{4}"